Find the general solution to ry + 3x y + 5y=0 Rys
Answer #1

Hiyo of T find general Solution 22 yil +30 y + 5y = 0 that The equation we assume form of the in has a solution yox 8-) y =8(8-1) i +38x² + snozo x* (ver-1) +3r+s) - EO g²_873875 co . ar?+bx+(20 +2=-b£ 56²490 2 ४ + 28+5=o 2a here a=1, 6=2,c=s r=-2